OpenStudy (anonymous):

(9m) to the power -2

OpenStudy (cwrw238):

x to power to -2 = 1 / x^2 so its 1/ (9m)^2

OpenStudy (anonymous):

oh, so how would you answer it? or is that the answer?

OpenStudy (anonymous):

How did you get that up arrow?

OpenStudy (cwrw238):

thats a symbol for 'to the power'

OpenStudy (cwrw238):

(9m)^2 = 9^2 * m^2 = 81m^2

OpenStudy (cwrw238):

so answer is 1 / (81m^2)

OpenStudy (anonymous):

Oh okay, that is what I thought

OpenStudy (cwrw238):

ok - yw